Calculate the ΔH for the following reaction: F2(g) + Br2(g) → 2FBr(g)

Bond Type Bond Energy (kJ/mol)

F−F 159

Br−Br 193

F−Br 212
Answer #1

For given reaction

Hrxn = n*H(bonds broken) - n*H ( bonds formed)  

Where n are the number of moles

Hrxn = [1*H(F2) + 1*H(Br2) ] - ( 2*H (F-Br) ) = [(1mol*159 kJ/mol) + (1 mol * 193 kJ/mol) -(2 mol *212 kJ/mol) =159 kJ + 193 kJ - 424 kJ = -72 kJ